The sum of 3 consecutive EVEN integers is 174. What is the first integer?
Answer by unlockmath(1122)

Hello,
Let's set up an equation first. We'll let x be the first even number.
Then we can say x+2 would be the next even number, Right? Well, the next even number would be x+4. So let's add all those together like this:
x+x+2+x+4=174
We're almost there. This equation could be written as:
3x + 6 = 174
Subtract 6 from both sides gives us:

3x = 168
Divide both sides by 3 gives us:
x = 56
Therefore the first number is 56.
Check it out. Does 56 + 58 + 60 = 174
Yep it does.
You're welcome,
RJ Toftness
www.math-unlock.com